I AM HAVING TROUBLE SETTING UP THIS PROBLEM: "FIND THREE CONSECUTIVE ODD INTEGERS SUCH THAT THE SUM OF THE SMALLEST AND 7 TIMES THE LARGEST IS 68". I SET IT UP x+7x=68. WHAT DID I DO WRONG??
Answer by rfer(16322)

x=5
x+2=7
x+4=9
--------------
x+7(x+4)=68
x+7x+28=68
8x=40
x=5
---
7*9+5=68
You did not give a the proper value to the largest number.